# An interpretation of scalars in SO(32)

### v2

@@ Abstract
keywords:
supersymmetry , flavour
journal: Physics Letters B
[…]
[first]organization=Institute for Biocomputation and Physics of Complex Systems (BIFI), Universidad de Zaragoza,city=Zaragoza, postcode=E-50009, country=Spain
1 Introduction
While highly relevant in string theory and supergravity, the $SO(32)$ group is not a good unification group as it doesn’t have complex representations [18]. But it still gets an interesting family group when decomposed. In this letter, we first review the decomposition, interpret it as a group symmetry on scalars that could be supersymmetry partners of the Standard Model fermions, and then we present an interesting reconstruction of such scalars as composites. Besides, the interpretation has an uniqueness that limits the number of generations for the SM group.
@@ 2 The flavour group in $SO(32)$
We can arrive to the same result by chaining some branchings. A straighforward way is $SO(32)\supset SU(16)\times U(1)$,
@@
3 $SO(32)$ from postulates
Once we know that our aim is to get not the fermions but just the scalar partners of a Susy Standard Model, we can wonder if there is some set of postulates that isolates directly the flavour group, or at least the number of generations it has. It turns up, there is an amusing set of requirements that force this result.
@@ 3.1 Turtles and elephants
3.2 Colourless and coloured flavour groups
@@ 4 A case for one generation
The $SU(5)$ model was frequently used as foundational model in the peak of composite theories in the early eighties. In most cases the components in the $5$ were new particles, but some models considered to keep right fermion as elementary and only left as composite, and even there was some proposal [13] were the $5$ had a shared mix of preons and known fermions222 Literature reviews also mention ideas where only the leptons are composite but I have not found a concrete reference.
5 Discussion on related groups
5.1 On $SU(15)$
@@
For the group decomposition, similar results could be obtained with only $SO(30)$ or $SU(15)$ as coloured flavour group, or $SO(10)$ or $SU(5)$ as colourless flavour groups, or even with $Usp(32)$.
@@ 5.2 On $SU(8)$
Family GUT unification with $SU(8)$ was examined with some detail in 1980, see for instance the references in the recent revisit of [3]. Typically three families of standard model fermions were expected to be in the summed complex representation $\bar{8}\otimes\bar{2}8\otimes 56$ and some criteria was used to select the hypercharge assignments.Most models preferred to interpret for flavour the first $SU(3)$ in $SU(8)\supset SU(5)\otimes SU(3)\otimes U(1)$ instead of leaving it for colour as [17]. Both approaches differ only in the algebra of $U(1)$ charges for the multiplets. The fundamental decomposes as a colour triplet, a $SU(2)$ horizontal doublet, and a $SU(3)$ horizontal triplet.
@@ 6 Discussion
The focus on scalars, and thus in pairs of fermions, makes the results to differ from preon constructions with the same groups. To get again a fermion, one should consider an extra object, particle or string, providing an 1/2 spin.
7 Conclusions
In conclusion, lets review what we have got. We offer a novel interpretation of the $SO(32)$ group within the context of supersymmetric models, emphasizing its potential as a flavour group. The decomposition and hypercharge assignment that allows to recover three generations has not been presented in the literature explicitly. This is for the obvious reason that it recovers scalars, not fermions. But on the other hand, to look for scalars avoids to address the problem of the lack of chiral fermions on SO(32).
@@ 7 Conclusions
Our postulate is, certainly, exotic: it suggests that while SSM fermions could be -or not- elementary, the SSM scalars are composites, with their preons being a subset of the observed SM fermions. Far fetched as this postulate looks, it reproduces the $SO(32)$ decomposition and fixes the number of possible generations.
The decomposition seems to imply that each generation also includes two extra ”scalar quarks” of charge ±$\pm 4/3$. It is unclear if such scalars could have an associated fermion, as it should be of Weyl type, not Dirac.
8 Acknowledgements

### v3

@@ Abstract
I Introduction
While highly relevant in string theory and supergravity, the $SO(32)$ group is not a good unification group as it doesn’t have complex representations [19]. But it still gets an interesting family group when decomposed. In this letter, we first review the decomposition, interpret it as a group symmetry on scalars that could be supersymmetry partners of the Standard Model fermions, and then we present an interesting reconstruction of such scalars as composites. Besides, the interpretation has a uniqueness that limits the number of generations for the SM group.
@@ II The flavour group in $SO(32)$
We can arrive to the same result by chaining some branchings. A straightforward way is $SO(32)\supset SU(16)\times U(1)$,
@@
III $SO(32)$ from postulates
Once we know that our aim is to get not the fermions but just the scalar partners of a Susy Standard Model, we can wonder if there is some set of postulates that isolates directly the flavour group, or at least the number of generations it has. It turns out, there is an amusing set of requirements that force this result.
@@ III.1 Turtles and elephants
III.2 Colourless and coloured flavour groups
@@ IV A case for one generation
The $SU(5)$ model was frequently used as foundational model in the peak of composite theories in the early eighties. In most cases the components in the $5$ were new particles, but some models considered to keep right fermion as elementary and only left as composite, and even there was some proposal [14] were the $5$ had a shared mix of preons and known fermions. The author in [4] considers a fundamental ”quint” and a composite $10$, albeit from the product of three ”quints”. Early literature also includes ideas where only the leptons are composite, as well as proposals where only the first family is elementary.
V The role of the top
The argument in section III tells us that there is one quark that does not act as a preon for the susy scalars. It does not tell us which one. We need to identify the elephant. And really we have not a concrete argument.
We favour the top quark due to the horizontal, flavour-like, symmetries found in the first section: $SU(3)_{f_{3}}\times SU(2)_{f_{2}}$. When considering the values of quark Yukawa couplings, and thus quark masses, it seems more fitting that the $f_{2}$ symmetry relates $u$ and $c$, with the $t$ quark being the excluded one. This scenario is typical when breaking textures.
We were in fact inspired by the empirical observation that toponium doesn’t exist, but this is perfectly justified by the mass of the top being closer to the Fermi scale than to the QCD scale. So it disintegrates faster than it binds. However, we have not discussed the binding mechanism here, and the one from QCD doesn’t apply.
The heavy mass of the top quark can be used as an argument in compositions where the quarks act as charges at the ends of a relativistic open string, as they need to be massless, or at least as light as possible. And ultimately it can be expected that masses in the standard model -with right neutrinos but not other particles- are protected by two dual symmetries: one that fixes all the degrees of freedom as Dirac massless except the top quark, and another one that fixes all the degrees of freedom as Majorana massless except the neutrinos  [^2] Such protection splits degrees of freedom as 84+12 in both cases, and thus we should look for some group representations in the $84$, $42$ or $21$. Note M2-brane and M5-brane carry a $SU(9)$ symmetry .
VI Discussion on Masses
The main point of this section is the existence of a mass formula that produces pairs of equal mass particles, as required by unbroken susy.
Excepting the top quark, preon models from the eighties were known to produce realistic masses. One model from that age is [24], that became popular later due to the accuracy in the lepton sector. It assumes that all the mass comes from a single abelian charge and that all the preons are presented in the same state, so the energy of a pair, having the same spatial wavefunction, is simply the energy of a single element with the sum of the charges
$E(q_{a},q_{b})=E(q_{a}+q_{b})=(q_{a}+q_{b})^{2}K_{\Omega}$ (14)
Intuitively one could imagine classical charge on two spherical surfaces of radius $\Lambda$ and common center, and sum both self-energies plus the interaction energy and see how it does not depend of the radius $\Lambda$.
Furthermore, the model [24] makes two provisions for composites of three pairs $(a,b^{i})$ to produce realistic masses:
  • That the charge of the three $b^{i}$ particles add to zero
    $z_{1}+z_{2}+z_{3}=0$ (15)
  • That the self energy of the common preon $a$ averages the self energy of the other three
    $z_{0}^{2}={z_{1}^{2}+z_{2}^{2}+z_{3}^{2}\over 3}$ (16)
The first condition is easily met extracting an abelian charge from any direction of a $SU(3)$ triplet, and we have some. The second is a sort of trace condition but it is imposed ad-hoc. Remember also that $z_{i}^{2}=(z_{j}+z_{k})^{2}$. So at least some scalars keep having the same self-energy than a third preon in the set.
Lets parametrize with an angle $\alpha$ all the possible ways to produce an abelian charge from the $3$ representation of SU(3), in $T_{3},T_{8}$ basis
$z_{1}=$ $\frac{1}{2}\cos\alpha$ $+\frac{1}{2\sqrt{3}}\sin\alpha$ (17)
$z_{2}=$ 00 0 $-\frac{1}{\sqrt{3}}\sin\alpha$ (18)
$z_{3}=$ $-\frac{1}{2}\cos\alpha$ $+\frac{1}{2\sqrt{3}}\sin\alpha$ (19)
$z_{0}=$$\pm 1/\sqrt{6}$ (20)
This is $T_{3}$ for $\alpha=0$ and $T_{8}$ for $\alpha=\pi/2$
The solutions have an obvious periodicity $32\pi/$ 3 and symmetries at $\pi/6$ and $\pi/2$.
As it is well known, if we use a scale factor $k=m_{e}+m_{\mu}+m_{\tau}$ then for $\alpha=0.745821$ the mass triplet $[k(z_{0}+z_{i})^{2}]$ recovers exactly the values $m_{e},m_{\mu},m_{\tau}$. This is sometimes interpreted as a prediction for $m_{\tau}$ [24], as we can use $m_{e}$ and $m_{\mu}$ to recover $\alpha,k$ and then calculate the extant mass, well within one sigma of the current measurement error. Note also that $kz_{0}^{2}=313.85$ MeV, a familiar quantity from QCD [27].
More interestingly, we can ask for the mass values of the octet and sextet. Note that they do not depend on $z_{0}$; our choosing of sign is translated to the whole $[z_{1},z_{2},z_{3}]$ tuple, but this is in turn just a exchange of preons with antipreons.
0 $\pi/2$ $\pi/4$ $\alpha_{SM}$ $\alpha_{.68583}$
$\bar{u}d$ 1553.4 914.63 1756.96 1776.9 1801.22
(3,2) $\bar{u}s$ 313.85 53.848 0 0.5110 3.4179
$\bar{u}b$ 15.853 914.63 126.144 105.65 78.463
$(\bar{3},2)$ (as above)
$\bar{s}d,\bar{d}s$ 470.775 1412.32 1756.96 1717.2 1647.7
(8,1) $\bar{d}b,\bar{b}d$ 470.775 1412.32 126.144 91.47 49.128
$\bar{b}s,\bar{s}b$ 1883.1 0 941.55 1016.0 1127.8
$(\bar{s}s,\bar{d}d,\bar{b}b)$ 0
$\bar{u}u,\bar{c}c$ 0
$(\bar{u}c,\bar{c}u)$ 0 if $u=c$, 1255.4 if $u\neq c$
Table 1: ”slepton” masses for $kz_{0}^{2}=$313.85 MeV. There are always two massless sneutrinos in the octet (8,1) and other two in (1,3). The other two combinations of (1,3) and (1,1) can be chosen to have zero mass too.
0 $\pi/2$ $\pi/4$ $\alpha_{SM}$ $\alpha_{.68583}$
$ud$ 15.853 26.93 42.048 45.18 49.128
(3,2) $us$ 313.85 1829.25 1255.4 1205.3 1127.8
$ub$ 1553.4 26.93 585.65 632.66 706.162
$dd$ 1883.1 627.7 2342.61 2388.9 2445.29
$ds$ 470.775 156.93 42.048 55.31 78.463
(6,1) $ss$ 0 2510.8 1255.4 1156.1 1007.05
$bd$ 0 627.7 313.85 289.03 251.76
$bs$ 470.775 156.93 585.65 597.21 611.32
$bb$ 1883.1 627.7 168.192 221.23 313.85
Table 2: ”squark” masses for $kz_{0}^{2}=$313.85 MeV. Note that a global change of sign in $s,b,c$ exchanges the $(3,2)$ of sleptons and squarks
VI.1 Paired scalars in the same representation
Our goal is to check if we can realistically recover a mass spectrum similar to supersymmetry scalars. This means that we should find pairs of scalars in the same representation having equal masses. We have to cases.
For $\alpha=\pi/2$ or $\pi/6$ we recover two equal masses and a different one in the triplet of $(3,2)$
$\frac{1}{4}\pm\frac{\sqrt{2}}{6},\frac{1}{2}\mp\frac{\sqrt{2}}{3},\frac{1}{4}\pm\frac{\sqrt{2}}{6}$
The octet is never a problem as we have the ”antiparticles” in the same set. The sextet fails for one pair. The triplets of $(3,2)$ are also a source of trouble, as the pair should be expected to happen across the $SU(2)$; the only possibility is to assign the same charge to ”preons” $c$ and $u$.
A better result happens for $\alpha=0$. Here a change of sign just exchanges charges $z_{1}$ and $z_{3}$, so we get exact pairs of masses in all cases. More important, also the sextet is grouped into pairs.
So we can conclude that it is possible, with this mass formula, to obtain an spectrum that resembles the scalars of a mildly broken supersymmetry.
If we take seriously that the mass of the fermions is just the ”preonic” self energy, then the discrepancy in the $\alpha=0$ case contains a sort of isospin exchange: we have assigned
$\begin{gathered}u=c=313.85,\\ s=0,d=b=470.8\end{gathered}$
and we have got
$\begin{gathered}(u,c,t)=(0,470.8,1883),\\ (d,s,b)=(15.8,313.85,1553.4)\end{gathered}$
VI.2 Solutions with paired scalars in different representation. Missing symmetry.
Lets go a bit beyond this the scope of this work, to look also for rotations that create some pairs in different representations. The motivation here is to explore for some extra symmetry that includes quarks of different charges. This is phenomenologically motivated by two approximate observations:
  • that the top mass still can fit in a Koide SU(3) tuple but only respect to high masses, for instance $(m_{c},m_{b},m_{t})$.
  • that $(m_{s},m_{c},m_{b})$ also look as a valid tuple.
Both observations need to get values from different representations. Looking for such solutions is beyond the scope of this work, but we can cover here the cases where a mass appears in two representations, as it is really a completion of our inspection.
We find two cases. First, for $\alpha=\pi/4=0.7854...$ we obtain for the triplet of the sleptonic $(3,2)$:
$\frac{1}{4}(2+\sqrt{3}),0,\frac{1}{4}(2-\sqrt{3})$ (21)
and with change of sign:
$\frac{1}{12}(2-\sqrt{3}),\frac{2}{3},\frac{1}{12}(2+\sqrt{3})$ (22)
The second case happens when $z_{0}=\pm 2z_{i}$ for some index $i$. All the solutions are similar obtained as reflections and translations of $\sin^{-1}{1\over 2\sqrt{2}}$ ; we show in the tables $\alpha=0.6858...$ for continuity.
The proportion (21) was first found by [21] who assigned them to $(m_{s},m_{u},m_{d})$, back in 1978.
It is interesting that if we scale (21,22) to get equal masses in both tuples, we observe a relation $\sum m_{q}=3\sum m_{l}$, and that works empirically for $(m_{s},m_{c},m_{b})$ with respect to $(m_{e},m_{\mu},m_{\tau})$. Also a comparison of table 2 respect to the quantities of table 1 seems to hint a missed factor of three in other cases: $627.7\to 1883.1$, $156.93\to 470.7$, $313.85\to 841.55$. With the factor three the tuple we have called $(ud,us,ub)$ gets more realistic masses and we know from elsewhere that it can be further rotated to produce more exact values, but we will not pursue this here 333 It is just an empirical observation, that solving $m_{3}=\left(\left(\sqrt{m_{1}}+\sqrt{m_{2}}\right)\left(2-\sqrt{3+6\frac{\sqrt{m_{1}m_{2}}}{\left(\sqrt{m_{1}}+\sqrt{m_{2}}\right)^{2}}}\right)\right)^{2}$ with input (172.4, 4.183) gets 1.3495, and then input (4.183,1.3495) gets 0.092
Our conclusion in this section is that the breaking down to a single $SU(3)$ flavour, and particularly the neglect of weak isospin, misses some extra symmetry that would allow more comprehensive mass formulae.
VII Discussion on related groups
VII.1 On $SU(15)$
@@
For the group decomposition, similar results could be obtained with only $SO(30)$ or $SU(15)$ as a coloured flavour group, or $SO(10)$ or $SU(5)$ as colourless flavour groups, or even with ${Usp}(32)$.
@@ VII.2 On $SU(8)$
Family GUT unification with $SU(8)$ was examined with some detail in 1980, see for instance the references in the recent revisit of [3]. Typically three families of standard model fermions were expected to be in the summed complex representation $\bar{8}+\bar{2}8+56$ and some criteria was used to select the hypercharge assignments.Most models preferred to interpret for flavour the first $SU(3)$ in $SU(8)\supset SU(5)\otimes SU(3)\otimes U(1)$ instead of leaving it for colour as [18]. Both approaches differ only in the algebra of $U(1)$ charges for the multiplets. The fundamental decomposes as a colour triplet, a $SU(2)$ horizontal doublet, and a $SU(3)$ horizontal triplet.
@@ VIII Discussion
If we get the scalars from $SO(32)$ a natural question is where the superpartners -the actual fermions- are. They could be obtained by applying the susy transformation. In string-inspired GUTs, they should be in usual non-susy models for three generations, and the issue would be to reassure the compatibility with the $SO(32)$ group. In preon constructions, our focus on scalars, and thus in pairs of fermions, makes the results to differ from most previous approaches. To recover a fermion, one must consider an extra object, particle or string, providing again an 1/2 spin.
IX Conclusions
In conclusion, lets review what we have got. We offer a novel interpretation of the $SO(32)$ group within the context of supersymmetric models, emphasizing its potential as a flavour group for scalars. The decomposition and hypercharge assignment that allows to recover three generations has not been presented in the literature explicitly. This is for the obvious reason that it recovers scalars, not fermions. But on the other hand, to look for scalars avoids to address the problem of the lack of chiral fermions.
@@ IX Conclusions
Our postulate is, certainly, exotic: it suggests that while SSM fermions could be elementary, the SSM scalars are composites, with their preons being a subset of the observed SM fermions. Far fetched as this postulate looks, it reproduces the $SO(32)$ decomposition and fixes the number of possible generations. Also, it justifies the non observation of superpartners; supersymmetry could be hiding in plain sight, not broken but distorted.
We have recalled a classical mass formula that when applied to our piece of $SU(3)$ flavour symmetry can grant pairs of scalars having the same mass.
The decomposition seems to imply that each generation also includes two extra ”scalar quarks” of charge ±$\pm 4/3$. It is unclear if such scalars could have an associated fermion, as it should be of Weyl type, not Dirac. On other hand, our approach could be compared to the ”scalar democracy” of [22] that goes further and proposes the existence of a scalar bound state for every pair of fundamental fermions, either leptons or quarks.
X Acknowledgements